OpenStudy (anonymous):

One number is four plus one half of another. The sum of the numbers is 31. Find the numbers.

OpenStudy (freckles):

can you translate the sentences into mathematical equations?

OpenStudy (freckles):

close

OpenStudy (freckles):

one number is 4 plus half of another x=4+1/2*y

OpenStudy (anonymous):

Yea I can write an equation but I don't know what to do next

OpenStudy (alexandervonhumboldt2):

solve the system of equatons

OpenStudy (freckles):

plug x=4+y/2 into the second equation x+y=31 see you can replace x there with 4+y/2

OpenStudy (freckles):

then solve for y

OpenStudy (freckles):

then solve for x

OpenStudy (anonymous):

Omg thank you so much